Mr. Li is 4 times the age of his son. Write the sum of their ages 6 years ago.
Answer by Alan3354(69443)

Mr. Li is 4 times the age of his son. Write the sum of their ages 6 years ago.
---------
L = 4S
Sum = 4S + 4S = 8S
---
6 years ago:
Sum = L - 6 + 4S - 6 = L + 4S - 12
----------
Or,
Sum = 8S - 12